# Life-Code: Central Dogma Modeling with Multi-Omics Sequence Unification

### v1

@@ A.4 Supervised Fine-tuning
In most cases, we apply Supervised Fine-tuning (SFT) to transfer pre-trained models to downstream tasks. Following (Nguyen et al., 2024c; Zhou et al., 2024), adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy (Hu et al., 2021) is employed to parameter-efficiently fine-tuning the models by AdamW optimizer with a batch size of 32. For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate {1e-5, 5e-5, 1e-4}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench (Liu et al., 2024a). Note that the maximum input length will be determined for different tasks since the sequence lengths of downstream tasks vary widely. We report the averaged results over three runs with the optimal settings.
@@ B.1 DNA Tasks with Genomics Benchmark
As proposed by (Grešová et al., 2023), three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark. As for the enhancer prediction, three datasets are provided for identifying enhancer regions in the mouse or human genome. As for the species classification, two datasets are selected for identifying sequences as either coding (exonic) or intergenic (non-coding) and classifying sequences as originating from humans or worms (C. elegans). As for the regulatory elements classification, three datasets are used for classifying sequences as regulatory regions based on Ensembl annotations, identifying open chromatin regions, or identifying non-TATA promoter regions in the human genome. We utilize the fully reproduced results of various DNA models in GenBench (Liu et al., 2024a).
@@ B.2 DNA Tasks with GUE Benchmark
As proposed by DNABERT2 (Zhou et al., 2024), the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric. To comprehensively evaluate the genome foundation models in modeling variable-length sequences, tasks with input lengths ranging from 70 to 1000 are selected. The following descriptions of the supported tasks are included in the GUE benchmark, where these resources are attached for illustration.
Promoter Detection (Human). This task identifies human proximal promoter regions essential for transcription initiation. Accurate detection aids in understanding gene regulation and disease mechanisms. The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) (Dreos et al., 2013). Meanwhile, we construct the non-promoter class with equal-sized randomly selected sequences outside of promoter regions but with TATA motif (TATA non-promoters) or randomly substituted sequences (non-TATA, non-promoters). We also combine the TATA and non-TATA datasets to obtain a combined dataset named all.
@@
Transcription Factor Binding Site Prediction (Human). This task predicts human transcription factor binding sites (TFBS), crucial for gene expression regulation. Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) (Consortium et al., 2012b) are collected via the UCSC genome browser. TFBS sequences are 101-bp regions around peaks, while non-TFBS sequences match in length and GC content. There are 5 datasets selected from a curated subset of 690, excluding trivial or overly challenging tasks.
Splice Site Prediction (Human). This task predicts splice donor and acceptor sites, the exact locations in the human genome where alternative splicing occurs. This prediction is crucial to understanding protein diversity and the implications of aberrant splicing in genetic disorders. The dataset (Scalzitti et al., 2021) consists of 400-bp-long sequences extracted from Ensembl GRCh38 human reference genome. As suggested by Ji et al. (2021a), existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality. As such, we reconstruct the dataset by iteratively adding adversarial examples (unseen false positive predictions in the hold-out set) in order to make this task more challenging.
Transcription Factor Binding Site Prediction (Mouse). This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) (Consortium et al., 2012a) from the UCSC genome browser. Negative examples are created by di-nucleotide shuffling. Five datasets are randomly selected from the 78 datasets using the same process as the human TFBS prediction dataset.
@@
Covid Variant Prediction (Virus). This task aims to predict the variant type of the SARS_CoV_2 virus based on 1000-length genome sequences. We download the genomes from the EpiCoV database (Khare et al., 2021) of the Global Initiative on Sharing Avian Influenza Data (GISAID). We consider 9 types of SARS_CoV_2 variants, including Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda and Zeta.
@@
Following (Zhou et al., 2024; Shen et al., 2024), we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression. During splicing, introns are removed from precursor messenger RNAs (pre-mRNAs), and exons are joined together to form mature mRNAs. This process is essential for generating functional mRNAs that could be translated into proteins. Identifying splice sites—the donor sites at the 5’ end of introns and the acceptor sites at the 3’ end—is vital for accurately predicting gene structure and location. Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset (Wang et al., 2019b) used in the GUE benchmark.
Spliceator dataset. This dataset (Scalzitti et al., 2021) consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans. The gold standard dataset GS 1 is adopted, which contains an equal number of positive and negative samples, and the F1 score is used as the evaluation metric. We chose three independent test datasets containing the samples from 3 different species of humans, fish (Danio rerio), and fruit fly (Drosophila melanogaster).
SpliceAI dataset. This dataset (Jaganathan et al., 2019) also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information. SpliceAI also references the ClinVar database to evaluate the clinical significance of predicted splicing variants, which contains information on clinically relevant variants and their associations with diseases. This dataset can be regarded as a long-range evaluation and adopts the top-1 AUC-ROC score as the metric.
@@ B.4 Protein and Multi-omic Tasks
Following EVO (Nguyen et al., 2024a) and protein language models (Lin et al., 2023b), we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness (Notin et al., 2022). EVO obtained (DMS) datasets with bacterial (prokaryote) and human (eukaryote) proteins from ProteinGYM at https://proteingym.org. To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies (Notin et al., 2022; Altae-Tran et al., 2021). For generative pre-trained models such as EVO, we rely on likelihood-based scores under the same masking scheme, assessing how well the model anticipates mutations. The model performances of zero-shot function prediction are measured by the strength of Spearman’s Rank Correlation Coefficient (SRCC) that correlates the predicted likelihoods with the experimental fitness measurements. Full results of protein fitness prediction are shown in Table 5 and Table A6, in which our Life-Code achieves balancing performances with bacterial and human proteins and surpasses NT-2500M and EVO-7B.
@@
Following LucaOne (He et al., 2024), we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins. Since experimentally identifying ncRPI) is typically expensive and time-consuming, the AI-based ncRPI can be a promising task. LucaOne proposes a binary classification task involving pairs of ncRNA and Amino Acid sequences (20,824 pairs in total) with top-1 accuracy as the metric.
@@
To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne (He et al., 2024) to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated. LucaOne collects a total of 8,533 accurate DNA-protein pairs from 13 species in the NCBI RefSeq database, with each DNA sequence extending to include an additional 100 nucleotides at both the 5’ and 3’ contexts, preserving intron sequences within the data. LucaOne generated double the number of negative samples by implementing substitutions, insertions, and deletions within the DNA sequences or altering amino acids in the protein sequences to ensure the resultant DNA sequences could not be accurately translated into their respective proteins.

### v2

@@ A.4 Supervised Fine-tuning
In most cases, we apply Supervised Fine-tuning (SFT) to transfer pre-trained models to downstream tasks. Following [36, 64], adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy [15] is employed to parameter-efficiently fine-tuning the models by AdamW optimizer with a batch size of 32. For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate $1\times 10^{-5}$, $5\times 10^{-5}$, $1\times 10^{-4}$}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench [28]. Note that the maximum input length will be determined for different tasks since the sequence lengths of downstream tasks vary widely. We report the averaged results over three runs with the optimal settings.
@@ B.1 DNA Tasks with Genomics Benchmark
As proposed by [12], three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark. As for the enhancer prediction, three datasets are provided for identifying enhancer regions in the mouse or human genome. As for the species classification, two datasets are selected for identifying sequences as either coding (exonic) or intergenic (non-coding) and classifying sequences as originating from humans or worms (C. elegans). As for the regulatory elements classification, three datasets are used for classifying sequences as regulatory regions based on Ensembl annotations, identifying open chromatin regions, or identifying non-TATA promoter regions in the human genome. We utilize the fully reproduced results of various DNA models in GenBench [28].
@@ B.2 DNA Tasks with GUE Benchmark
As proposed by DNABERT2 [64], the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric. To comprehensively evaluate the genome foundation models in modeling variable-length sequences, tasks with input lengths ranging from 70 to 1000 are selected. The following descriptions of the supported tasks are included in the GUE benchmark, where these resources are attached for illustration.
Promoter Detection (Human). This task identifies human proximal promoter regions essential for transcription initiation. Accurate detection aids in understanding gene regulation and disease mechanisms. The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) [10]. Meanwhile, we construct the non-promoter class with equal-sized randomly selected sequences outside of promoter regions but with TATA motif (TATA non-promoters) or randomly substituted sequences (non-TATA, non-promoters). We also combine the TATA and non-TATA datasets to obtain a combined dataset named all.
@@
Transcription Factor Binding Site Prediction (Human). This task predicts human transcription factor binding sites (TFBS), crucial for gene expression regulation. Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) [6] are collected via the UCSC genome browser. TFBS sequences are 101-bp regions around peaks, while non-TFBS sequences match in length and GC content. There are 5 datasets selected from a curated subset of 690, excluding trivial or overly challenging tasks.
Splice Site Prediction (Human). This task predicts splice donor and acceptor sites, the exact locations in the human genome where alternative splicing occurs. This prediction is crucial to understanding protein diversity and the implications of aberrant splicing in genetic disorders. The dataset [46] consists of 400-bp-long sequences extracted from Ensembl GRCh38 human reference genome. As suggested by [17], existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality. As such, we reconstruct the dataset by iteratively adding adversarial examples (unseen false positive predictions in the hold-out set) in order to make this task more challenging.
Transcription Factor Binding Site Prediction (Mouse). This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) [7] from the UCSC genome browser. Negative examples are created by di-nucleotide shuffling. Five datasets are randomly selected from the 78 datasets using the same process as the human TFBS prediction dataset.
@@
Covid Variant Prediction (Virus). This task aims to predict the variant type of the SARS_CoV_2 virus based on 1000-length genome sequences. We download the genomes from the EpiCoV database [21] of the Global Initiative on Sharing Avian Influenza Data (GISAID). We consider 9 types of SARS_CoV_2 variants, including Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda and Zeta.
@@
Following [64, 49], we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression. During splicing, introns are removed from precursor messenger RNAs (pre-mRNAs), and exons are joined together to form mature mRNAs. This process is essential for generating functional mRNAs that could be translated into proteins. Identifying splice sites—the donor sites at the 5’ end of introns and the acceptor sites at the 3’ end—is vital for accurately predicting gene structure and location. Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset [56] used in the GUE benchmark.
Spliceator dataset. This dataset [46] consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans. The gold standard dataset GS 1 is adopted, which contains an equal number of positive and negative samples, and the F1 score is used as the evaluation metric. We chose three independent test datasets containing the samples from 3 different species of humans, fish (Danio rerio), and fruit fly (Drosophila melanogaster).
SpliceAI dataset. This dataset [16] also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information. SpliceAI also references the ClinVar database to evaluate the clinical significance of predicted splicing variants, which contains information on clinically relevant variants and their associations with diseases. This dataset can be regarded as a long-range evaluation and adopts the top-1 AUC-ROC score as the metric.
@@ B.4 Protein and Multi-omic Tasks
Following EVO [35] and protein language models [27], we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness [38]. EVO obtained (DMS) datasets with bacterial (prokaryote) and human (eukaryote) proteins from ProteinGYM at https://proteingym.org. To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies [38, 2]. For generative pre-trained models such as EVO, we rely on likelihood-based scores under the same masking scheme, assessing how well the model anticipates mutations. The model performances of zero-shot function prediction are measured by the strength of Spearman’s Rank Correlation Coefficient (SRCC) that correlates the predicted likelihoods with the experimental fitness measurements. Full results of protein fitness prediction are shown in Table 3 and Table A6, in which our Life-Code achieves balancing performances with bacterial and human proteins and surpasses NT-2500M and EVO-7B.
@@
Following LucaOne [13], we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins. Since experimentally identifying ncRPI) is typically expensive and time-consuming, the AI-based ncRPI can be a promising task. LucaOne proposes a binary classification task involving pairs of ncRNA and Amino Acid sequences (20,824 pairs in total) with top-1 accuracy as the metric.
@@
To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne [13] to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated. LucaOne collects a total of 8,533 accurate DNA-protein pairs from 13 species in the NCBI RefSeq database, with each DNA sequence extending to include an additional 100 nucleotides at both the 5’ and 3’ contexts, preserving intron sequences within the data. LucaOne generated double the number of negative samples by implementing substitutions, insertions, and deletions within the DNA sequences or altering amino acids in the protein sequences to ensure the resultant DNA sequences could not be accurately translated into their respective proteins.